# SEKE: Specialised Experts for Keyword Extraction

### v1

@@ Abstract
Keyword extraction involves identifying the most descriptive words in a document, allowing automatic categorisation and summarisation of large quantities of diverse textual data. Relying on the insight that real-world keyword detection often requires handling of diverse content, we propose a novel supervised keyword extraction approach based on the mixture of experts (MoE) technique. MoE uses a learnable routing sub-network to direct information to specialised experts, allowing them to specialize in distinct regions of the input space. SEKE, a mixture of Specialised Experts for supervised Keyword Extraction, uses DeBERTa as the backbone model and builds on the MoE framework, where experts attend to each token, by integrating it with a recurrent neural network (RNN), to allow successful extraction even on smaller corpora, where specialisation is harder due to lack of training data. The MoE framework also provides an insight into inner workings of individual experts, enhancing the explainability of the approach. We benchmark SEKE on multiple English datasets, achieving state-of-the-art performance compared to strong supervised and unsupervised baselines. Our analysis reveals that depending on data size and type, experts specialize in distinct syntactic and semantic components, such as punctuation, stopwords, parts-of-speech, or named entities. Code is available at https://github.com/matejMartinc/SEKE_keyword_extraction.
@@ 1 Introduction
Keyword extraction is crucial for efficiently summarizing, indexing, and retrieving relevant information from large textual corpora, making it one of the most fundamental NLP tasks Song et al. (2023). Automatically extracting keywords is defined as the automatic extraction of words that represent crucial semantic aspects of the text. While the first automated solutions have been proposed more than two decades ago (Tumey, 1999; Mihalcea and Tarau, 2004), the task is far from solved, and the scientific community is still actively trying to improve the performance of the keyword extractors, proposing several new algorithms in recent years (see Section 2 for details).
@@
Recently, the so-called foundation models have completely revolutionised the way natural language processing (NLP) is done, and the development of large language models (LLMs) has enabled task-agnostic architectures to solve downstream NLP tasks in a zero-shot fashion without the need for labelled data Brown et al. (2020). These breakthroughs have been enabled by the integration of novel training regimes and architectural components, such as Low-Rank Adaptation (LoRA) Hu et al. (2021) and Mixture of Experts (MoE) Jacobs et al. (1991), which make models more efficient, while requiring less computational resources.
While these components have been successfully employed in an unsupervised language model setting with abundant data, there are less studies exploring their effectiveness in a supervised setting with much fewer resources. The main reason for this is that these components were developed with a specific objective of improving the scalability of the models Shazeer et al. (2017) and therefore their effect on the performance of the models is somewhat neglected or measured only indirectly, that is, studies focus on how the use of these components allows for greater scalability, which in turn leads to performance gains.
@@
In contrast, in this study, we focus on the performance improvements obtained from these components without increasing the size of the backbone model. More specifically, we are interested in exploring the question of whether MoE can improve the performance of the model on a specific downstream task, in our case keyword extraction. The general hypothesis we want to test is whether it is beneficial for keyword extraction to allow specific parts of the network to specialize for specific types of tokens. More specifically, we explore if introducing a gating network that assigns different parts of the sequence to specialised layers, which would only attend to specific subsequences and tokens according to their semantic and grammatical role, could improve the current state-of-the-art (SOTA) in the field. In addition, we explore what type of specialisation different experts undertake and how much data is required for a successful specialisation. For these reasons, we test our approach on several datasets with different amounts of training data available, to determine the data threshold that still allows the convergence of expert layers. Finally, we examine whether synergy can be achieved between MoE and recurrent neural network (RNN) that would lower this threshold.
The contributions of the paper are as follows.
@@
  • We propose a novel token classification head architecture (see Figure 1) that combines MoE and RNN. The approach outperforms SOTA-supervised sequence-labelling models on most keyword extraction datasets. As far as we know, this is the first research that studies the usage of MoE in a supervised sequence-labelling scenario.
@@
  • We demonstrate that experts specialize in different syntactic and semantic structures, ranging from punctuation characters and part-of-speech tags to specific words and named entities, depending on the domain they are trained on and the size of the training data.
@@
  • We show that adding MoE layers during fine-tuning generally has a positive effect on the model’s performance no matter the dataset, but is somewhat dependent on the amount of data available. In addition, we show that the data requirements can be lowered by introducing additional recurrent layers.
@@ 2.1 Mixture of Experts
The idea for Mixture of experts (MoE) comes from the study by Jacobs et al. (1991). Like ensemble methods, the concept involved a supervised procedure for a system composed of separate networks, each responsible for a different subset of training cases. Each network, or expert, specializes in a distinct region of the input space. The choice of expert is managed by a gating network, which determines the weights for each expert. Both the experts and the gating network are trained during the training process.
More recently, the study by Eigen et al. (2013) proposed integrating MoE as components within deep neural networks. This allowed MoE to function as layers in a multilayer network, enabling models to be both large and efficient, which led to the employment of the MoE method in the NLP setting. The study by Shazeer et al. (2017) proposed a 137-billion-parameter LSTM model (the prevalent NLP architecture for language modelling at the time) by introducing sparsity, which allowed for very fast inference even at large scales. This work, which focused on translation, encountered several challenges such as high communication costs and training instabilities. After the success of the Mixtral 8x7B Jiang et al. (2024), MoE have been used in several LLMs because they enable more efficient pre-training with fewer computing resources, allowing greater scalability in terms of model size and data set size. The studies show that a MoE model achieves the same quality as its dense counterpart much faster during pre-training, leading to its usage in training multi-trillion parameter models, such as the 1.6T parameters Switch Transformers Fedus et al. (2022).
@@ 2.2 Keyword Extraction
Contemporary studies on keyword extraction treat it either as a text generation or a sequence labelling task. Yuan et al. (2020) proposed an encoder-decoder RNN architecture featuring two mechanisms, semantic coverage and orthogonal regularisation, which treated keyword extraction as text generation. These mechanisms ensure that the overall inner representation of a generated keyword sequence is semantically aligned with the overall meaning of the source text. In contrast, several recent sequence labelling approaches tackle the task with transformers. Sahrawat et al. (2020) tested several transformer architectures, namely BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and GPT-2, and two distinct sequence labelling heads, a bidirectional Long short-term memory head (BiLSTM) and a BiLSTM head with an additional Conditional random fields layer (BiLSTM-CRF). The TNT-KID approach Martinc et al. (2022) on the other hand adapted the transformer architecture specifically for keyword extraction tasks by re-parameterizing the attention mechanism to focus more on positional information. Finally, Sun et al. (2021) proposed JointKPE, an open domain keyphrase extraction (KPE) architecture also based on pre-trained transformer language models. It ranks keyphrases by assessing their informativeness across the entire document and is simultaneously trained on a keyphrase chunking task to ensure the phraseness of the keyphrase candidates.
@@
Just recently, Luo et al. (2023) proposed Diff-KPE, a text diffusion approach to generate enhanced keyphrase representations utilizing the supervised Variational Information Bottleneck (VIB). Diff-KPE generates keyphrase embeddings conditioned on the entire document using BERT and then injects the generated keyphrase embeddings into each phrase representation. By optimizing the ranking network and VIB simultaneously, they manage to rank each candidate phrase by utilizing both the information of keyphrases and the entire document.
Keyword extraction can also be tackled in an unsupervised manner. In general, these approaches can be divided into four main categories, namely statistical, graph-based, embeddings-based, and language model-based methods. Statistical methods, such as YAKE Campos et al. (2020), use the statistical characteristics of texts to capture keywords. Graph-based methods, such as TextRank Mihalcea and Tarau (2004) and RaKUn Škrlj et al. (2019) build graphs to rank words based on their position in the graph. Embedding-based methods, such as Key2Vec Mahata et al. (2018), employ semantic information from distributed word and sentence representations for keyword extraction.
@@
Notably, large language model-based methods tend to be highly effective, driven by the rise of LLMs, such as ChatGPT Brown et al. (2020). These models have been employed in several recent studies Martínez-Cruz et al. (2023); Luo et al. (2023), demonstrating promising results due to their zero-shot capabilities, which allow them to perform tasks without task-specific fine-tuning Kojima et al. (2022). However, Zhang et al. (2024) has recently highlighted the theoretical drawbacks of autoregressively trained LLM models for semantic classification tasks, attributing these limitations to their fixed token positioning, which restricts inter-sample connections. In contrast, masked language models, such as BERT Devlin et al. (2019), leverage flexible token targeting to overcome these issues. Building on these findings, we select the DeBERTa He et al. (2020) model as a strong backbone.
@@ 3.1 Architecture
Our approach (visualised in Figure 1) is based on fine-tuning a pre-trained transformer architecture, in our case DeBERTa He et al. (2020), with a specialised token classification head. The architecture was chosen since it showcased a SOTA performance on several downstream NLP tasks, among them also sequence-labelling tasks such as NER Shon et al. (2022). Additionally, we hypothesised that the model is especially appropriate for keyword extraction due to its novel disentangled attention mechanism. More specifically, in the DeBERTa model, each word is represented using two vectors that encode its content and position separately, and the attention weights among words are computed using disentangled matrices on their contents and relative positions, respectively. We assumed that this would allow the model to better distinguish between the positional and semantic/lexical information and therefore make it possible to assign attention to some tokens purely based on their position in the text, improving the overall performance. This hypothesis is based on previous study by Martinc et al. (2022) showing that token position is especially important in the keyword identification task due to the disproportionate distribution of keywords in the document, which is skewed towards the beginning of the document (i.e., more keywords appear at the beginning of the document). By considering the importance of positional information, they fed the positional encoding to the attention mechanism directly, subsequently managing to improve the performance.
The main novelty of our approach is a customised token classification head containing several components. The output logits from the DeBERTa backbone are first fed to the gate network or router that determines which tokens in the input sequence are sent to which expert. More specifically, in the setting with n experts, the gate network (G) decides which experts (E) receive a part of the input:
$y=\sum_{i=1}^{n}G(x)_{i}E_{i}(x)$
@@
The outputs of different experts, which attend to specific tokens assigned to them by the routing network, are first reassembled into a single sequence representation. Then, a two-layer randomly initialised encoder, consisting of dropout and two RNN layers, is added (with element-wise summation) to this sequence. The initial motivation behind this adaptation is connected with findings from the related work which suggest that recurrent layers are quite successful at modelling the positional importance of tokens in the keyword detection task and by the study of Martinc et al. (2022), who reported good results when an RNN classifier and contextual embeddings generated by transformers were employed for keyword detection.
[…]
We conduct additional post-processing, namely, the identified keyphrases are extracted from the labelled sequence produced by the model, and we conduct additional filtering, where we remove keyphrases containing punctuation (except dashes and apostrophes), and deduplication, where we remove duplicate keyphrases. The final output list contains a maximum of 10 keyphrases per input document.
@@ 4.1 Datasets
We conduct experiments on selected benchmark datasets used in the SOTA papers to which we compare our approach. More specifically, we experiment on 6 datasets from the related work Meng et al. (2017); Martinc et al. (2022); Xiong et al. (2019), covering three domains, news, scientific articles, and web pages.
@@ 4.2 Experiments
We test different settings to determine the effect that the proposed architectural additions have on the performance of the model. We are interested in the specific contributions of the MoE and RNN layers to the performance of the model (either separately or together); therefore, we compare the settings containing these components to the baseline DeBERTa model with just a (usual) feed-forward token classification head.
We employ the DeBERTa model [^1] More specifically, the “deberta-v3-base” version available at https://huggingface.co/microsoft/deberta-v3-base. and fine-tune it using Low-Rank Adaptation (LoRA) Hu et al. (2021). We freeze all layers in the backbone model and train only low-rank perturbations to query and value weight matrices in the model. Additionally, we train all the matrices in the custom token classification head (i.e., MoE, RNN, and dense classification layers are randomly initialised and fine-tuned). We use the same hyperparameter setting across all datasets, namely, the following values were used during fine-tuning: learning rate of 2e-4, LoRA intrinsic rank of 16, LoRA alpha of 16, LoRA dropout of 0.1, and sequence length of 256 for all tasks.
In order to determine how does training dataset size effect our approach, we employ two distinct training scenarios. In the validation set training scenario, the DeBERTa model was fine-tuned just on each dataset’s validation sets, which were randomly split into 80 percent of documents used for fine-tuning and 20 percent of documents used for hyperparameter optimisation and test set model selection. We fine-tune the models for up to 20 epochs and employ early stopping. Note that the JPTimes dataset with no available predefined validation-test splits is only used for testing. Here, the model was fine-tuned on the KPTimes-valid dataset.
@@
To assess the performance of our models and compare the results of our keyword extraction approach to other SOTA methods, we adopt the same evaluation methodology as in Meng et al. (2017); Martinc et al. (2022); Luo et al. (2023). We use the same test splits for the sake of comparability between our approach and others, and measure F1@$k$ with $k\in\{1,3,5\}$ on the openKP dataset and $k\in\{5,10\}$ on other datasets. To determine the exact match, we first lowercase the candidate keywords and the gold standard keywords, and then apply Porter Stemmer Porter (1980) to both.
@@
DeBERTa RNN 31.2±1.0 31.5±1.0 47.7±0.7 56.6±0.8 29.2±1.5 29.6±1.5 44.1±0.7 44.2±0.7 36.2±1.2 36.6±1.3 26.8±1.1 37.2±0.5 38.1±0.4
DeBERTa MoE RNN 32.7±1.1 33.0±1.1 47.2±1.1 56.5±1.0 29.5±1.1 30.1±1.0 44.9±0.3 45.1±0.3 37.4±0.5 37.8±0.6 27.1±0.9 38.9±0.5 40.1±0.4
@@
DeBERTa RNN 34.4±0.6 34.7±0.6 47.6±0.6 57.5±0.8 32.3±1.5 32.6±1.4 53.5±0.1 53.5±0.1 38.1±0.9 38.2±1.0 31.9±2.3 39.8±1.0 40.2±0.8
DeBERTa MoE RNN 36.0±0.3 36.5±0.3 49.4±1.0 58.8±0.4 34.3±0.9 34.7±0.9 54.8±1.0 55.0±0.9 39.0±1.7 39.3±1.8 33.3±1.5 42.5±0.6 42.9±0.5
Table 2: Empirical evaluation of the keyword extractors. We report the average over five runs (with five different random seeds) and the standard deviation. Bold represents the best score according to a specific criterion on a specific dataset.

### v2

@@ Abstract
Keyword extraction involves identifying the most descriptive words in a document, allowing automatic categorisation and summarisation of large quantities of diverse textual data. Relying on the insight that real-world keyword detection often requires handling of diverse content, we propose a novel supervised keyword extraction approach based on the mixture of experts (MoE) technique. MoE uses a learnable routing sub-network to direct information to specialised experts, allowing them to specialise in distinct regions of the input space. SEKE, a mixture of Specialised Experts for supervised Keyword Extraction, uses DeBERTa as the backbone model and builds on the MoE framework, where experts attend to each token, by integrating it with a bidirectional Long short-term memory (BiLSTM) network, to allow successful extraction even on smaller corpora, where specialisation is harder due to lack of training data. The MoE framework also provides an insight into inner workings of individual experts, enhancing the explainability of the approach. We benchmark SEKE on multiple English datasets, achieving state-of-the-art performance compared to strong supervised and unsupervised baselines. Our analysis reveals that depending on data size and type, experts specialise in distinct syntactic and semantic components, such as punctuation, stopwords, parts-of-speech, or named entities. Code is available at https://github.com/matejMartinc/SEKE_keyword_extraction.
SEKE: Specialised Experts for Keyword Extraction
Matej Martinc  and Hanh Thi Hong Tran  and Senja Pollak  and Boshko Koloski Jožef Stefan Institute, Ljubljana, Slovenia {name}.{surname}@ijs.si
@@ 1 Introduction
Keyword extraction, i.e., extraction of words that represent crucial semantic aspects of the text, is crucial for efficiently summarizing, indexing, and retrieving relevant information from large textual corpora, making it one of the most fundamental NLP tasks Song et al. (2023). While the first automated solutions have been proposed more than two decades ago (Tumey, 1999; Mihalcea and Tarau, 2004; Hulth, 2003), the task is far from solved, and the scientific community is still actively trying to improve the performance of the keyword extractors (see Section 2 for details).
@@
Recently, the so-called foundation models have completely revolutionised the way natural language processing (NLP) is done, and the development of large language models (LLMs) has enabled task-agnostic architectures to solve downstream NLP tasks in a zero-shot fashion without the need for labelled data (Brown et al., 2020). While recent related studies (Martínez-Cruz et al., 2024; Luo et al., 2024) and experiments conducted in this study show that LLMs’ keyword extraction performance is still not comparable to specialised supervised approaches, their zero-shot performance nevertheless remains impressive. It was made possible, to a large extent, by the integration of novel training regimes and architectural components, such as Low-Rank Adaptation (LoRA) Hu et al. (2021) and Mixture of Experts (MoE) Jacobs et al. (1991), which make models more efficient, while requiring fewer computational resources. While these components have been successfully employed in an unsupervised language modelling setting with abundant data, there are less studies exploring their effectiveness in a supervised setting with much fewer resources. The main reason for this is that these components were developed with a specific objective of improving the scalability of the models Shazeer et al. (2017) and therefore their effect on the performance is somewhat neglected or measured only indirectly, that is, studies focus on how the use of these components allows for greater scalability, which in turn leads to performance gains.
@@
In contrast, in this study, we focus on the performance improvements obtained from these components without increasing the size of the backbone model. The hypothesis we want to test is whether it is beneficial for keyword extraction to allow specific parts of the network to specialise for specific types of tokens. Since the “keywordiness” of a token is to a large extent determined by its specific contextual role, we explore if introducing a gating network that assigns different parts of the sequence to specialised layers, which would only attend to specific subsequences and tokens according to their semantic and grammatical role, could improve the current state-of-the-art (SOTA) in the field. In addition, we explore what type of specialisation different experts undertake and how much data is required for a successful specialisation. We test our approach on several datasets with different amounts of training data available, to determine the data threshold that still allows the convergence of expert layers. Finally, we examine whether synergy can be achieved between MoE and bidirectional Long short-term memory (BiLSTM) network that would lower this threshold and improve overall performance by amplifying the sequential information lost by the context slicing caused by the MoE strategy. The contributions of the paper are as follows.
@@
  • We propose a novel token classification head architecture (see Figure 1) that combines MoE and BiLSTM. The approach outperforms other SOTA keyword extraction approaches on most datasets. As far as we know, this is the first research that studies the usage of MoE in a supervised sequence-labelling scenario.
@@
  • We demonstrate that experts specialise in different syntactic and semantic structures, ranging from punctuation characters and part-of-speech tags to specific words and named entities, depending on the domain they are trained on and the size of the training data.
@@
  • We show that adding MoE layers during fine-tuning generally has a positive effect on the model’s performance no matter the dataset, but is somewhat dependent on the amount of data available. In addition, we show that the data requirements can be lowered by introducing additional BiLSTM layers.
@@ 2.1 Mixture of Experts
The idea for Mixture of experts (MoE) comes from the study by Jacobs et al. (1991). The concept involved a supervised procedure for a system composed of separate networks, each responsible for a different subset of training cases. Each network, or expert, specialises in a distinct region of the input space. The choice of expert is managed by a gating network, which determines the weights for each expert. Both the experts and the gating network are trained during the training process.
More recently, the study by Eigen et al. (2013) proposed integrating MoE as components within deep neural networks. This allowed MoE to function as layers in a multilayer network, enabling models to be both large and efficient, which led to the employment of the MoE method in the NLP setting. The study by Shazeer et al. (2017) proposed a 137-billion-parameter LSTM model (the prevalent NLP architecture for language modelling at the time) by introducing sparsity, which allowed for very fast inference even at large scales. This work, which focused on translation, encountered several challenges such as high communication costs and training instabilities. After the success of the Mixtral 8x7B Jiang et al. (2024), MoE have been used in several LLMs because they enable more efficient pre-training with fewer computing resources, allowing greater scalability in terms of model and data set size. The studies show that a MoE model achieves the same quality as its dense counterpart much faster during pre-training, leading to its usage in training multi-trillion parameter models Fedus et al. (2022).
@@ 2.2 Keyword Extraction
Contemporary studies on keyword extraction treat it either as a text generation or a sequence labelling task. Yuan et al. (2020) proposed an encoder-decoder RNN architecture featuring two mechanisms, semantic coverage and orthogonal regularisation, which treated keyword extraction as text generation. These mechanisms ensure that the representation of a generated keyword sequence is semantically aligned with the overall meaning of the source text. In contrast, several recent sequence labelling approaches tackle the task with transformers. Sahrawat et al. (2020) tested several transformer architectures, namely BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and GPT-2, and two distinct sequence labelling heads, a BiLSTM and a BiLSTM head with an additional Conditional random fields layer (BiLSTM-CRF). The TNT-KID approach Martinc et al. (2022) on the other hand adapted the transformer architecture specifically for keyword extraction tasks by re-parameterizing the attention mechanism to focus more on positional information. Finally, Sun et al. (2021) proposed JointKPE, an open domain keyphrase extraction (KPE) architecture also based on pre-trained transformer language models. It ranks keyphrases by assessing their informativeness across the entire document and is simultaneously trained on a keyphrase chunking task to ensure the phraseness of the keyphrase candidates.
@@
Just recently, Luo et al. (2024) proposed Diff-KPE, a text diffusion approach to generate enhanced keyphrase representations utilizing the supervised Variational Information Bottleneck (VIB). Diff-KPE generates keyphrase embeddings conditioned on the entire document using BERT and then injects the generated keyphrase embeddings into each phrase representation. Another recent approach is HybridMatch Song et al. (2024), which consists of the representation-focused Siamese encoder component and the interaction-based component that estimates relatedness between candidate phrases and the corresponding document.
@@
Keyword extraction can also be tackled in an unsupervised manner. In general, these approaches can be divided into four main categories, namely statistical (e.g., YAKE Campos et al. (2020)), graph-based (e.g., TextRank Mihalcea and Tarau (2004) and RaKUn Škrlj et al. (2019)), embeddings-based (e.g., Key2Vec Mahata et al. (2018)), and language model-based methods. Notably, large language model-based methods tend to be highly effective, driven by the rise of LLMs, such as ChatGPT Brown et al. (2020). These models have been employed in several recent keyword extraction studies Martínez-Cruz et al. (2024); Luo et al. (2024); Lee et al. (2023); Peng et al. (2024), demonstrating promising results due to their zero-shot capabilities, which allow them to perform tasks without task-specific fine-tuning Kojima et al. (2022). However, while significantly outperforming other unsupervised approaches, these models still do not offer performance comparable to the SOTA supervised methods. To some extent, this might be related to the deficiency observed by Zhang et al. (2024), which has highlighted the theoretical drawbacks of autoregressively trained LLM models for semantic classification tasks, attributing these limitations to their fixed token positioning, which restricts inter-sample connections. In contrast, masked language models, such as BERT Devlin et al. (2019) and DeBERTa He et al. (2020), leverage flexible token targeting to overcome these issues. Additionally, through fine-tuning they can be easily (and cheaply) adapted to each specific keyword assignment regime, which vary across different domains.
@@ 3.1 Architecture
Our approach (visualised in Figure 1) is based on fine-tuning a pre-trained transformer architecture, in our case DeBERTa He et al. (2020), with a specialised token classification head. The architecture was chosen since it showcased a SOTA performance on several downstream NLP tasks, among them also sequence-labelling tasks such as NER Shon et al. (2022). Additionally, we hypothesised that the model is especially appropriate for keyword extraction due to its novel disentangled attention mechanism. In the DeBERTa model, each word is represented using two vectors that encode its content and position separately, and the attention weights among words are computed using disentangled matrices on their contents and relative positions, respectively. We assumed that this would allow the model to better distinguish between the positional and semantic/lexical information and therefore make it possible to assign attention to some tokens purely based on their position in the text, improving the overall performance. This hypothesis is based on a previous study by Martinc et al. (2022) showing that token position is especially important in the keyword identification task due to the disproportionate distribution of keywords in the document, which is skewed towards the beginning of the document (i.e., more keywords appear at the beginning of the document). By considering the importance of positional information, they fed the positional encoding to the attention mechanism directly, subsequently managing to improve the performance.
The main novelty of our approach is a customised token classification head with MoE. It is motivated by the hypothesis that token-based specialization offered by MoE might be especially useful for keyword extraction. The specialization of different attention heads and attention layers in the transformer architecture has been extensively analysed in the past, also in the context of keyword extraction (Martinc et al., 2022). This specialization is mostly embeddings-based, i.e., different parts of the embedding space for each token are specialized. In contrast, the specialization enabled by MoE occurs on the sequence level, since the router decides for each token in the sequence to which expert it should go. Since the “keywordiness” of a specific token (or a sequence of tokens) to a large extent depends on a context, we assumed that this type of specialization might bring benefits.
The output logits from the DeBERTa backbone are first fed to the gate network or router that determines which tokens in the input sequence are sent to which expert. More specifically, in the setting with n experts, the gate network (G) decides which experts (E) receive a part of the input:
[…]
@@
The outputs of different experts, which attend to specific tokens assigned to them by the routing network, are first reassembled into a single sequence representation. Then, a two-layer randomly initialised encoder, consisting of dropout and two BiLSTM layers, is added (with element-wise summation) to this sequence. The employment of the BiLSTM module is there to amplify the sequential information lost by the context slicing caused by the MoE strategy, i.e., the weakening of the positional information due to overspecialization. This is especially problematic in case of multi-word keyphrases, where different parts of the keyphrase are attended by different experts. This adaptation is also motivated by the findings from the related work Sahrawat et al. (2020), where it was shown that putting a BiLSTM classification head on top of different transformer models improved the keyword extraction capabilities of the model.
The output sequence of the MoE and BiLSTM encoders is finally fed to the feed-forward classification layer, which returns the output matrix of size $\textrm{SL}*\textrm{NC}$, where SL stands for sequence length and NC stands for the number of classes. In our case, the $NC$ is 3, since we model the keyword extraction task as sequence labelling and employ the so-called BIO labelling approach (same as, for example, in Sahrawat et al. (2020)). Using the BIO annotation regime, each word in the document is assigned one of the three possible labels: $k_{b}$ denotes that the word is the first word in a keyphrase, $k_{i}$ means that the word is inside a keyphrase, and $k_{o}$ indicates that the word is not part of a keyphrase.
The identified keyphrases are extracted from the labelled sequence produced by the model, and we conduct additional filtering, where we remove keyphrases containing punctuation (except dashes and apostrophes), and deduplication, where we remove duplicate keyphrases. The final output list contains a maximum of 10 keyphrases per input document.
@@ 4.1 Datasets
We conduct experiments on selected benchmark datasets used in the SOTA papers to which we compare our approach. More specifically, we experiment on 6 publicly available datasets with MIT license from the related work Meng et al. (2017); Martinc et al. (2022); Xiong et al. (2019), covering three domains, news, scientific articles, and web pages.
@@ 4.2 Experiments
We test different settings to determine the effect that the proposed architectural additions have on the performance of the model. We are interested in the specific contributions of the MoE and BiLSTM layers to the performance of the model (either separately or together); therefore, we compare the settings containing these components to the baseline DeBERTa model with just a (usual) feed-forward token classification head.
We employ the DeBERTa model [^1] More specifically, the “deberta-v3-base” version available at https://huggingface.co/microsoft/deberta-v3-base under the MIT license. and fine-tune it using Low-Rank Adaptation (LoRA) Hu et al. (2021). We freeze all layers in the backbone model and train only low-rank perturbations to query and value weight matrices in the model. Additionally, we train all the matrices in the custom token classification head (i.e., MoE, BiLSTM, and dense classification layers are randomly initialised and fine-tuned). See Section A for details about the hyperparameter setting used.
In order to determine how the size of the training dataset affects our approach, we employ two distinct training scenarios. In the validation set training scenario, the DeBERTa model was fine-tuned just on each dataset’s validation sets, which were randomly split into 80 percent of documents used for fine-tuning and 20 percent of documents used for hyperparameter optimisation and test set model selection. We fine-tune the models for up to 20 epochs and employ early stopping. Note that the JPTimes dataset with no available predefined validation-test splits is only used for testing. Here, the model was fine-tuned on the KPTimes-valid dataset.
@@
To assess the performance of our models and compare the results of our keyword extraction approach to other SOTA methods, we adopt the same evaluation methodology as in Meng et al. (2017); Martinc et al. (2022); Luo et al. (2024). We use the same test splits for the sake of comparability between our approach and others, and measure F1@$k$ with $k\in\{1,3,5\}$ on the openKP dataset and $k\in\{5,10\}$ on other datasets [^2] Note that the values of $k$ differ between openKP and other datasets. This is due to the variability of keyword assignment regimes across different domains, i.e., news and scientific articles on average have more keywords than web pages (see Table 1 for details). We report the same F1@$k$ for each dataset as in related work to facilitate easy comparison.. To determine the exact match, we first lowercase the candidate keywords and the gold standard keywords, and then apply Porter Stemmer Porter (1980) to both.
@@
RaKUN 17.7 16.0 10.1 10.8 12.7 10.6 16.8 13.9 22.5 18.5 8.81 10.12 6.45
YAKE 14.1 14.6 20.4 22.3 21.5 19.6 10.5 11.8 10.9 13.5 3.73 4.24 2.94
Large Language Model baselines
ChatGLM2-6b / / 25.1 30.1 / / / / / / 16.0 11.0 8.6
@@
DeBERTa BiLSTM 31.2±1.0 31.5±1.0 47.7±0.7 56.6±0.8 29.2±1.5 29.6±1.5 44.1±0.7 44.2±0.7 36.2±1.2 36.6±1.3 26.8±1.1 37.2±0.5 38.1±0.4
DeBERTa MoE BiLSTM 32.7±1.1 33.0±1.1 47.2±1.1 56.5±1.0 29.5±1.1 30.1±1.0 44.9±0.3 45.1±0.3 37.4±0.5 37.8±0.6 27.1±0.9 38.9±0.5 40.1±0.4
@@
DeBERTa BiLSTM 34.4±0.6 34.7±0.6 47.6±0.6 57.5±0.8 32.3±1.5 32.6±1.4 53.5±0.1 53.5±0.1 38.1±0.9 38.2±1.0 31.9±2.3 39.8±1.0 40.2±0.8
DeBERTa MoE BiLSTM 36.0±0.3 36.5±0.3 49.4±1.0 58.8±0.4 34.3±0.9 34.7±0.9 54.8±1.0 55.0±0.9 39.0±1.7 39.3±1.8 33.3±1.5 42.5±0.6 42.9±0.5
Table 2: Empirical evaluation of the keyword extractors. We report the average over five runs and the standard deviation. Bold represents the best score according to a specific criterion on a specific dataset.